A Bibliography of IEC Standards for Computing Systems and Programming Languages

Nelson H. F. Beebe Center for Scientific Computing

University of Utah

Department of Mathematics, 322 INSCC

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254

FAX: +1 801 585 1640, +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@ieee.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

24 April 2000 Version 1.00

Title word cross-reference

10967-1 [IEC94a]. 1539-1 [IEC97]. 1539-2 [IEC94b]. 1539-3 [IEC99]. 1981-01 [IEC81]. 1985-12 [IEC85]. 1988-12 [IEC88]. 1989-01 [IEC89]. 1990-12 [IEC90]. 1994-12 [IEC94a, IEC94b]. 1997-12 [IEC97]. 1998-12 [IEC98a, IEC98b]. 1999-02 [IEC99].

8651-1 [IEC88].

9593-1 [IEC90].

Application [IEC85]. arithmetic [IEC89, IEC94a].

Base [IEC97]. Binary [IEC89]. bindings [IEC88, IEC90].

CAMAC [IEC81]. character [IEC94b]. compilation [IEC99]. Computer [IEC88, IEC90]. Conditional [IEC99]. control [IEC85].

data [IEC98b].

Enhanced [IEC98b]. exception [IEC98a].

facilities [IEC98b]. Floating [IEC98a, IEC89, IEC94a]. Floating-point [IEC98a, IEC89]. FORTRAN [IEC85, IEC88, IEC90, IEC94b, IEC97, IEC98a, IEC98b, IEC99].

GKS [IEC88]. Graphical [IEC88]. Graphics [IEC88, IEC90].

handling [IEC98a]. Hierarchical [IEC90].

independent [IEC94a]. Industrial [IEC85]. Information [IEC88, IEC90, IEC94a, IEC94b, IEC97, IEC98a, IEC98b, IEC99]. Integer [IEC94a]. Interactive [IEC90]. ISO [IEC85, IEC88, IEC90, IEC94a, IEC94b, IEC97, IEC98a, IEC98b, IEC99]. ISO/IEC [IEC85, IEC88, IEC90, IEC94a, IEC94b, IEC97, IEC99]. ISO/IEC/TR2 [IEC98a, IEC98b].

Kernel [IEC88].

Language [IEC94a, IEC88, IEC90, IEC97]. languages [IEC94b, IEC97, IEC98a, IEC98b, IEC99]. length [IEC94b].

microprocessor [IEC89].

Part [IEC88, IEC90, IEC94a, IEC94b, IEC97, IEC99]. PHIGS [IEC90]. point [IEC89, IEC94a, IEC98a]. processes [IEC85]. processing [IEC88, IEC90]. Programmer [IEC90]. Programming [IEC94b, IEC97, IEC98a, IEC98b, IEC99].

real [IEC85]. real-time [IEC85].

strings [IEC94b]. Subroutines [IEC81]. System [IEC88, IEC90]. systems [IEC88, IEC89, IEC90].

technology [IEC94a, IEC94b, IEC97, IEC98a, IEC98b, IEC99]. time [IEC85]. TR2 [IEC98a, IEC98b]. type [IEC98b].

Varying [IEC94b].

References

IEC:1981:ISC

[IEC81] IEC. IEC 60713 (1981-01): Subroutines for CAMAC. International Electrotechnical Commission, 3, rue de Varembé, PO Box 131, CH-1211 Geneva 20, Switzerland. Telephone: +41 22 919 02 11. Telefax: +41 22 919 03 00. E-mail: info@iec.ch.

                                                URL: http://www.iec.ch, 1981. 57

pp. US$98.00. URL http://www.

                                                iec.ch/cgi-bin/procgi.pl/www/
                                                iecwww.p?wwwlang=E&wwwprog=cat-
                                                det.p&wartnum=017001.

IEC:1985:III

[IEC85] IEC. ISO/IEC 7846 (1985-12): Industrial real-time FORTRAN — Application for the control of industrial processes. International Electrotechnical Commission, 3, rue de Varembé, PO Box 131, CH-1211 Geneva 20, Switzerland. Telephone: +41 22 919 02 11. Telefax: +41 22 919 03 00. E-mail: info@iec.ch.

                                                URL: http://www.iec.ch, 1985. 32

pp. US$63.00. URL http://www.

                                                iec.ch/cgi-bin/procgi.pl/www/
                                                iecwww.p?wwwlang=E&wwwprog=cat-
                                                det.p&wartnum=018044.

IEC:1988:III

[IEC88] IEC. ISO/IEC 8651-1 (1988-12): Information processing systems — Computer Graphics — Graphical Kernel System (GKS) language bindings — Part 1: FOR-TRAN. International Electrotechnical Commission, 3, rue de Varembé, PO Box 131, CH-1211 Geneva 20, Switzerland. Telephone: +41 22 919 02 11. Telefax: +41 22 919 03 00. E-mail: info@iec.ch. URL:

http://www.iec.ch, 1988. 116

pp. US$156.00. URL http://www.

iec.ch/cgi-bin/procgi.pl/www/
iecwww.p?wwwlang=E&wwwprog=cat-
det.p&wartnum=018128.

IEC:1989:IBF

[IEC89] IEC. IEC 60559 (1989-01): Binary floating-point arithmetic for microprocessor systems. International Electrotechnical Commission, 3, rue de Varembé, PO Box 131, CH-1211 Geneva 20, Switzerland. Telephone: +41 22 919 02 11. Telefax: +41 22 919 03 00. E-mail: info@iec.ch.

URL: http://www.iec.ch, 1989. 45

pp. US$86.00. URL http://www.

iec.ch/cgi-bin/procgi.pl/www/
iecwww.p?wwwlang=E&wwwprog=cat-
det.p&wartnum=019113. This Stan-

dard was formerly known as IEEE 754.

IEC:1990:III

[IEC90] IEC. ISO/IEC 9593-1 (1990-12): Information processing systems — Computer graphics — Programmer’s Hierarchical Interactive Graphics System (PHIGS) language bindings — Part 1: FORTRAN. International Electrotechnical Commission, 3, rue de Varembé, PO Box 131, CH-1211 Geneva 20, Switzerland. Telephone: +41 22 919 02 11. Telefax: +41 22 919 03 00. E-

mail: info@iec.ch. URL: http:
//www.iec.ch, 1990.  213 pp.

US$230.00. URL http://www.

iec.ch/cgi-bin/procgi.pl/www/

                                        iecwww.p?wwwlang=E&wwwprog=cat-
                                        det.p&wartnum=018250.

IEC:1994:IIIa

[IEC94a] IEC. ISO/IEC 10967-1 (1994-12): Information technology — Language independent arithmetic — Part 1: Integer and floating point arithmetic. International Electrotechnical Commission, 3, rue de Varembé, PO Box 131, CH-1211 Geneva 20, Switzerland. Telephone: +41 22 919 02 11. Telefax: +41 22 919 03 00. E-mail: info@iec.ch. URL:

                                        http://www.iec.ch, 1994. 92 pp.

US$136.00. URL http://www.

                                        iec.ch/cgi-bin/procgi.pl/www/
                                        iecwww.p?wwwlang=E&wwwprog=cat-
                                        det.p&wartnum=019729.

IEC:1994:IIIb

[IEC94b] IEC. ISO/IEC 1539-2 (1994-12) Information technology — Programming languages — Fortran — Part 2: Varying length character strings. International Electrotechnical Commission, 3, rue de Varembé, PO Box 131, CH-1211 Geneva 20, Switzerland. Telephone: +41 22 919 02 11. Telefax: +41 22 919 03 00. E-mail: info@iec.ch. URL:

                                        http://www.iec.ch, 1994. 66 pp.

US$110.00. URL http://www.

                                        iec.ch/cgi-bin/procgi.pl/www/
                                        iecwww.p?wwwlang=E&wwwprog=cat-
                                        det.p&wartnum=017866.

IEC:1997:III

[IEC97] IEC. ISO/IEC 1539-1 (1997-12): Information technology — Programming languages — Fortran — Part 1: Base language. International Electrotechnical Commission, 3, rue de Varembé, PO Box 131, CH-1211 Geneva 20, Switzerland. Telephone: +41 22 919 02 11. Telefax: +41 22 919 03 00. E-

mail: info@iec.ch. URL: http:
//www.iec.ch, 1997.  346 pp.

US$249.00. URL http://www.

iec.ch/cgi-bin/procgi.pl/www/
iecwww.p?wwwlang=E&wwwprog=cat-
det.p&wartnum=022616.

IEC:1998:IITa

[IEC98a] IEC. ISO/IEC/TR2 15580 (1998-12): Information technology — Programming languages — Fortran — Floating-point exception handling. International Electrotechnical Commission, 3, rue de Varembé, PO Box 131, CH-1211 Geneva 20, Switzerland. Telephone: +41 22 919 02 11. Telefax: +41 22 919 03 00. E-mail: info@iec.ch.

URL: http://www.iec.ch, 1998. 27

pp. US$56.00. URL http://www.

iec.ch/cgi-bin/procgi.pl/www/
iecwww.p?wwwlang=E&wwwprog=cat-
det.p&wartnum=023454.

IEC:1998:IITb

[IEC98b] IEC. ISO/IEC/TR2 15581 (1998-12): Information technology — Programming languages — Fortran — Enhanced data type facilities. International Electrotechnical Commission, 3, rue de Varembé, PO Box 131, CH-1211 Geneva 20, Switzerland. Telephone: +41 22 919 02 11. Telefax: +41 22 919 03 00. E-mail: info@iec.ch.

URL: http://www.iec.ch, 1998. 13

pp. US$33.00. URL http://www.

iec.ch/cgi-bin/procgi.pl/www/

                                        iecwww.p?wwwlang=E&wwwprog=cat-
                                        det.p&wartnum=023455.

IEC:1999:III

[IEC99] IEC. ISO/IEC 1539-3 (1999-02): Information technology — Programming languages — Fortran — Part 3: Conditional compilation. International Electrotechnical Commission, 3, rue de Varembé, PO Box 131, CH-1211 Geneva 20, Switzerland. Telephone: +41 22 919 02 11. Telefax: +41 22 919 03 00. E-mail: info@iec.ch. URL:

                                        http://www.iec.ch, 1999.  21

pp. US$47.00. URL http://www.

                                        iec.ch/cgi-bin/procgi.pl/www/
                                        iecwww.p?wwwlang=E&wwwprog=cat-
                                        det.p&wartnum=023629.